Find the total capacitance for three capacitors connected in series, given their


individualcapacitances are 2.000nF, 5.000pF, and 8.000 μF.

1
Expert's answer
2022-04-04T17:09:26-0400
"\frac{1}{C}=\frac{1}{C_1}+\frac{1}{C_2}+\frac{1}{C_3}"

"\frac{1}{C}=\frac{1}{2.000*10^{-9}}+\frac{1}{5.000*10^{-12}}+\frac{1}{8.000*10^{-6}}"

"C=4.988*10^{-12}\:\rm F=4.988\: pF"